# CTC-DRO: Robust Optimization for Reducing Language Disparities in Speech Recognition

### v2

@@ Abstract
Modern deep learning models often achieve high overall performance, but consistently fail on specific subgroups. Group distributionally robust optimization (group DRO) addresses this problem by minimizing the worst-group loss, but it fails when group losses misrepresent performance differences between groups. This is common in domains like speech, where the widely used connectionist temporal classification (CTC) loss scales with input length and varies with linguistic and acoustic properties, leading to spurious differences between group losses. We present CTC-DRO, which addresses the shortcomings of the group DRO objective by smoothing the group weight update to prevent overemphasis on consistently high-loss groups, while using input length-matched batching to mitigate CTC’s scaling issues. We evaluate CTC-DRO on the task of multilingual automatic speech recognition (ASR) across five language sets from the ML-SUPERB 2.0 benchmark. CTC-DRO consistently outperforms group DRO and CTC-based baseline models, reducing the worst-language error by up to $47.1\%$ and the average error by up to $32.9\%$. CTC-DRO can be applied to ASR with minimal computational costs, and offers the potential for reducing group disparities in other domains with similar challenges.
Distributionally Robust Optimization, Deep learning, Robustness, Speech recognition
@@ 1 Introduction
Distributionally robust optimization (DRO), particularly group DRO (Sagawa et al., 2020), has the potential to mitigate language disparities in multilingual ASR. group DRO improves group robustness by dynamically up-weighting high-loss groups during training. However, it requires comparable training losses between groups to perform well (Oren et al., 2019; Sagawa et al., 2020), and this condition is often not met in ASR model training, because of differences in input length and speaker and acoustic characteristics across language-specific datasets. In this paper we focus on a training approach that has been successful in recent multilingual ASR benchmarks: pre-trained self-supervised models fine-tuned with the connectionist temporal classification (CTC; Graves et al., 2006) objective (Rouditchenko et al., 2023; Chen et al., 2024; Pratap et al., 2024). Differences in CTC-based training losses due to length, speaker, and acoustics may lead to varying magnitudes and irreducible components of losses across different groups. As a result, some groups with disproportionately high losses may dominate training with group DRO, causing under-training of the other groups, and ultimately negatively impacting overall downstream performance (Słowik & Bottou, 2022).
To address these issues, we present CTC-DRO, which optimizes a generalization of the group DRO objective, specifically by smoothing the up-weighting of high-loss groups. This new objective prevents overemphasis on groups with consistently and disproportionately high training losses. Also, we use length-matched group losses to mitigate the scaling properties of CTC. We evaluate CTC-DRO using language subsets randomly selected from the ML-SUPERB 2.0 (Shi et al., 2024) benchmark. In this setting, CTC-DRO models outperform both group DRO and CTC-based baseline models across five language sets. Specifically, CTC-DRO models reduce the error rate of the worst-performing language in all of the five sets, with improvements of up to $47.1\%$, while also improving the average performance across all languages by up to $32.9\%$. Our code and newly trained models are publicly available at https://github.com/Bartelds/ctc-dro.
@@ 2.1 Group DRO
Given a family of models $\Theta$, loss function $l$ and training data $(x,y)$ drawn from empirical distribution $\hat{P}$, the standard training procedure for label prediction involves minimizing the expected loss over the training data:

### v3

@@ Abstract
Modern deep learning models often achieve high overall performance, but consistently fail on specific subgroups. Group distributionally robust optimization (group DRO) addresses this problem by minimizing the worst-group loss, but it fails when group losses misrepresent performance differences between groups. This is common in domains like speech, where the widely used connectionist temporal classification (CTC) loss not only scales with input length but also varies with linguistic and acoustic properties, leading to spurious differences between group losses. We present CTC-DRO, which addresses the shortcomings of the group DRO objective by smoothing the group weight update to prevent overemphasis on consistently high-loss groups, while using input length-matched batching to mitigate CTC’s scaling issues. We evaluate CTC-DRO on the task of multilingual automatic speech recognition (ASR) across five language sets from the diverse ML-SUPERB 2.0 benchmark. CTC-DRO consistently outperforms group DRO and CTC-based baseline models, reducing the worst-language error by up to $47.1\%$ and the average error by up to $32.9\%$. CTC-DRO can be applied to ASR with minimal computational costs, and, while motivated by multilingual ASR, offers the potential for reducing group disparities in other domains with similar challenges.
@@ 1 Introduction
Distributionally robust optimization (DRO), particularly group DRO (Sagawa et al., 2020), has the potential to mitigate language disparities in multilingual ASR. Group DRO improves group robustness by up-weighting high-loss groups during training, and has been shown to outperform other approaches where the goal is to achieve high performance, even on the worst-performing group (Koh et al., 2021). However, it requires comparable training losses between groups to perform well (Oren et al., 2019; Sagawa et al., 2020), and this condition is often not met in ASR model training, because of differences in input length and speaker and acoustic characteristics across language-specific datasets.
In this paper, we focus on a training approach that has been successful on multilingual ASR benchmarks: pre-trained self-supervised models fine-tuned with the connectionist temporal classification (CTC; Graves et al., 2006) objective (Rouditchenko et al., 2023; Chen et al., 2024; Pratap et al., 2024). CTC-based models built on encoders such as XLS-R (Babu et al., 2022) and MMS (Pratap et al., 2024) are widely adopted and offer advantages over autoregressive models like Whisper  (Radford et al., 2023), including faster inference and reduced hallucinations (Koenecke et al., 2024; Peng et al., 2024), which are crucial for many downstream applications. However, differences in CTC-based training losses due to length, speaker, and acoustics may lead to varying magnitudes and irreducible components of losses across different groups. As a result, the group DRO weights do not have the desired effect.
To address these issues, we present CTC-DRO, which optimizes a generalization of the group DRO objective, specifically by smoothing the up-weighting of high-loss groups. This new objective prevents overemphasis on groups with consistently and disproportionately high training losses. Also, we use length-matched group losses to mitigate the scaling properties of CTC. We evaluate CTC-DRO using language sets randomly selected from the ML-SUPERB 2.0 (Shi et al., 2024) benchmark collection, which includes multilingual speech data from 15 diverse corpora across multiple domains, speaking styles and recording conditions. In this setting, CTC-DRO models outperform both group DRO and CTC-based baseline models across five language sets, regardless of whether balanced or unbalanced amounts of training data per language are used during training. Specifically, CTC-DRO models reduce the error rate of the worst-performing language in all of the five sets, with improvements of up to $47.1\%$, while also improving the average performance across all languages by up to $32.9\%$. While motivated by multilingual ASR, CTC-DRO offers the potential for reducing group disparities in other domains with incomparable training losses between groups, such as medical applications (Ganz et al., 2021; Petersen et al., 2023). Our code and newly trained models are publicly available at https://github.com/Bartelds/ctc-dro.
@@ 2.1 Group DRO
Given a family of models $\Theta$, loss function $\ell$ and training data $(x,y)$ drawn from empirical distribution $\hat{P}$, the standard training procedure for label prediction involves minimizing the expected loss over the training data: